OpenStudy (anonymous):

Can someone explain to me how to do this? Im confused.. 4 square root 6/ square root 30

OpenStudy (anonymous):

\[\frac{4\sqrt{6}}{\sqrt{30}}\]

OpenStudy (anonymous):

multiply top and bottom by \(\sqrt{30}\) to get \[\frac{4\sqrt{6}}{\sqrt{30}}\times \frac{\sqrt{30}}{\sqrt{30}}=\frac{4\sqrt{180}}{30}\]

OpenStudy (anonymous):

then we have more work to do

OpenStudy (anonymous):

since \(180=36\times 5\) the numerator is \[4\times \sqrt{180}=4\times \sqrt{36\times 5}=4\sqrt{36}\times \sqrt{5}=4\times 6\times \sqrt{5}=24\sqrt{5}\]

OpenStudy (anonymous):

so we are left with \[\frac{24\sqrt{5}}{30}\]and your last job is to cancel whatver you can

OpenStudy (anonymous):

but there is nothing to cancel out right? cause arnt you supposed to cancel out like terms?

OpenStudy (anonymous):

hello?